OpenStudy (anonymous):

solve x+y=5 3y+2x=11

OpenStudy (anonymous):

y=x-5 slope is 1

OpenStudy (anonymous):

DIVIDING equation 2 and substacting from 1 we get -y/2=-1/2 y=1 and x=4 hence solution is (4,1)